A square piece of cardboard is 12 inches on a side. A square piece x inches on a side is cut out from each corner. The flaps are then turned up to form an open box. Find polynomials that represent the volume and outside surface area of the box.

Answer by nyc_function(2741)

volume = w*h*l

volume = (12-2x)^3

volume = (x+y)(x+y)(x+y)

V = xxx+xxy+xyx+xyy+yxx+yxy+yyx+yyy

V = xxx+3xxy+3xyy+yyy

V = 12(12)(12)+3(12)(12)(-2x)+3(12)(-2x)(-2x…

V = 1728-864x+144x^2-8x^3

area = 5(12-2x)^2

area = 5(144-48x+4x^2)

area = 720-240x+20x^2